djil is 30 years older than diane. 5 years ago, djil's age is twice the age of diane one year from now . how old is djil now?
Answer by Kartik-The Math specialist(4)

Let the age of diane be x.
Djil's age=x+30
5 years before djil age was
x+30-5=x+25
1 year from now diane age will be x+1
ATQ
2(x+1)=x+25
x=23
Djil's age= x+30 = 23+30 = 53 years.